OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x helppppp

OpenStudy (anonymous):

Use the double angle formulas for sin and cos, then simplify.

OpenStudy (anonymous):

could you help me do that whats th first step?

OpenStudy (anonymous):

How about a couple of formulas?\[\cos 2x=\cos^2x-\sin^2x\]\[\sin 2x=2 \cos x \sin x\]

OpenStudy (anonymous):

2 cos x sin x - cos^2x-sin^2 x

OpenStudy (anonymous):

I think you have a sign error on the last term.